# thanks: These authors contributed equally to this work

### v1

@@ Abstract
The original proposal of quantum key distribution (QKD) was based on ideal single photon sources, which 40 years later, are still challenging to develop. Therefore, the development of decoy state protocols using weak coherent states (WCS) from lasers, set the frontier in terms of secure key rates. Here, we propose and experimentally emulate two simple-to-implement protocols that allow practical, far from ideal sub-Poissonian photon sources to outperform state-of-the-art WCS. By engineering the photon statistics of a quantum dot’s biexciton-exciton cascade, we show that either a truncated decoy state protocol or a heralded purification protocol can be employed to achieve a significantly increased performance in terms of the maximal allowed channel loss for secure key creation, which can exceed that of WCS by more than 3dB. We then show that our recently demonstrated room temperature single photon sources, based on giant colloidal quantum dots coupled to nano-antennas, are already well within the optimal performance range. These protocols can be utilized efficiently on a host of various sub-Poissonian quantum emitters having controllable photon statistics, offering a practical approach to QKD without the hindering requirements on the single photon purity of the photon source.
@@ I Introduction
Quantum key distribution (QKD) stands as a prominent candidate for the post-quantum-computer secure optical communication, enabling a secure key-establishing protocol between two parties employing the quantum nature of light [1, 2, 3]. In theory, encoding the information for QKD onto pure single photon states, is ultimately secured. In practice, it is limited by imperfections of the transmitter and the receiver, making security proofs a challenging task [4, 5], thus reducing the maximal secure key rate (SKR) and the maximal losses in the communication channel over which SKR is achievable [6, 7].
@@
Decoy state protocols introduce variations in the distribution of the emitted photons to better characterize the channel and obtain a tighter bound on the SKR [19, 20, 21, 22], allowing detection of PNS attacks [23, 24]. Yet, even with decoy state protocols, WCS are highly sensitive [25, 26] to the high probability of the vacuum state, due to the dependence between the different photon number emission probabilities, which overall limits the possible SKR [9]. The above solutions and inherent limitations of WCS for QKD set clear opportunities for improving the security and efficiency of QKD applications using realistic SPS systems [27].
@@
While efforts to develop a nearly ideal SPS are still ongoing by many groups [28, 29, 30, 31, 32], here we introduce a radically different, more realistic approach. By utilizing two new simple QKD protocols on compact, imperfect SPS sources, we significantly enhance the QKD performance, in terms of the MCL, as compared to even ideal, infinite decoy state protocols using WCS.
@@
We base our protocols on our recent demonstrations of room temperature operating SPS devices based on a giant colloidal nanocrystal quantum dot (gCQD) coupled to a hybrid nanoantenna and plasmonic nanocone [33, 34, 35, 36, 37]. These SPS devices displayed a highly enhanced rate of photon emission approaching the GHz range, together with near-unity collection efficiencies ($>90\%$) of the emitted photons. The photon emission from such gCQDs is based on the bi-exciton - exciton (BX-X) emission cascade under optical excitation [38], and has been shown to be well described by a truncated photon-number Fock space with $N=0,1$, and $2$ photons, with a negligible probability of $N>2$ photons [39].
@@
The two protocols are both based on our ability to vary and control the probabilities of emitting the different Fock states, $\{P_{0},P_{1},P_{2}\}$ over a wide range of values, by simply varying the optical excitation power. The first protocol we implement is a simple and realistic decoy state protocol for such a source. We show that with this new protocol, we can significantly exceed, by more than $3$ dB, the MCL value of an ideal decoy state protocol using a WCS source. The second protocol uses an alternative approach based on heralded photon purification under saturated excitation. We show that a similar performance enhancement can also be achieved. Finally we experimentally verify that the probability for $N>2$ ($P_{>2}$) is negligibly small, justifying our protocols’ assumptions, and demonstrating that even existing SPS which are far from being ideal, can outperform the current state-of-the-art protocols based on WCS sources.
@@
Figure 1: (a) Schematic of the BX-X cascade in an optically excited gCQD, using a non-resonant pulsed excitation with a normalized intensity $S=\frac{I}{I_{S}}$. Inset is a sketch of a typical gCQD, consisting of a CdSe core and a CdS shell. (b) Normalized emission counts (blue open circles) as a function of $S$, displaying a saturation behaviour. The dashed black line is a fit to a model, detailed in Section II. The orange circles show the experimentally extracted $g^{(2)}(0)$ for the different powers. (c) Photon probabilities as a function of $S$. The dashed black lines mark the intensities chosen for the signal $S_{2}$, and the two decoy states $S_{0}$, $S_{1}$, for the decoy state on a truncated basis (DTB) protocol.
@@ II Optical control of photon statistics from the BX-X emission cascade in a gCQD
Fig. 1a presents a schematic sketch of the BX-X cascade in a gCQD (see Supp. Sec. I) following a non-resonant pulse excitation with power $I$. Modelling this process, a BX state |$\ket{XX}$ is excited by the absorption of two pump photons. This process happens with a probability:
@@
Since $P_{XX},P_{X}$ depend on the excitation power, the probabilities to emit two, one or zero photons following excitation become excitation power dependent, allowing an external control over the emitted photon statistics as is required for decoy states protocols. Such a saturable behavior of the excitonic population allows us to define a relative excitation power $S=I/I_{S}$, where $I_{S}$ is the saturation power, for which the detected count rate (which is linearly proportional to the average emission photon number) reaches $90\%$ of the maximum value.
@@
In Fig. 1(b), we experimentally demonstrate photon statistics control by varying the excitation power from a single bare gCQD. As $S$ increases, the collected photon emission intensity increases sublinearly and saturates at $S\gtrsim 1$. A good fit of the theoretical model, based on the absorption probabilities given in Eqs. 1, 2 and on the subsequent emission probabilities $QY_{X},QY_{XX}$ is shown (see Supp. Sec. II). This fit confirms the basic assumptions given above for the BX-X cascade process.
@@
At the same time, the measured second-order correlation function, $g^{(2)}(0)$, also increases in a similar manner. This is understood in the following way: at low powers, most emission events are empty as $P_{0}$ is the dominant term. As $S$ increases, $P_{1}$ increases linearly (Eq. 2), while $P_{2}$ increases quadratically and thus is negligible at $S\ll 1$. Since $g^{(2)}(0)$ is an increasing function of $P_{2}/(\sum_{i=0,1,2}{P_{i}})$, a saturable behavior of $g^{(2)}(0)$ with increasing $S$ is also expected for $S\rightarrow 1$.
@@
Importantly, as we also show experimentally, multi-excitonic emissions from the gCQD can be neglected due to their low emission probability [39], thus $P_{N>2}=0$.
@@
By measuring both $g^{(2)}(0)$ and the count-rate $C$, we can extract the probability distribution precisely, as is detailed in Supp. Sec. III.
@@
In Fig. 1(c), we present the extracted photon emission probabilities $\{P_{0},P_{1},P_{2}\}$. As expected, the probability of emitting one or two photons increases with excitation power, but at a different rate, allowing optical control of the emitted photon statistics with only varying $S$.
@@
Figure 2: Concept of two protocols based on a high rate, high collection efficiency SPS with imperfect purity: (a) Experimental setup and sketch of the BB84+DTB protocol, with Alice adjusting the excitation power on the SPS between $S_{0}=0$, $S_{1}$ and $S_{2}$ to control the emitted photon statistics for the 2-decoy protocol, followed by collection optics with efficiency $\eta_{C}$ and a standard BB84 encoding unit. (b) Experimental setup and sketch of the BB84+HPP photon purification protocol. A single excitation power $S_{3}>1$, near saturation, is used to excite the SPS followed by collection optics with efficiency $\eta_{C}$ and a beam-splitter (BS) with a high efficiency single photon detector with detection efficiency $\eta_{D}$. Here only same pulse detections at Alice and Bob are used for the secure key.
@@ III.1 Decoy State on a Truncated Fock Basis - DTB
In Fig. 2(a), we present the conceptual experimental setup of the decoy state protocol. Here an SPS with a near unity photon collection efficiency, $\eta_{c}\simeq 1$ [36] (titled here SPS1), is used for implementing a BB84 protocol, but where Alice controls the excitation powers, $S_{0}=0$, $S_{1}$ and $S_{2}$. This in turn modifies the photon number statistics, thus allowing for the implementation of decoy and signal states [11, 15].
To establish the DTB protocol, we define the gain of the signal (decoyi), $Q_{s}$ ($Q_{d,i}$), as the fraction of encoded photon pulses sent by Alice and detected by Bob. We also define the error rate of the signal (decoyi), $E_{s}$ ($E_{d,i}$), as the fraction of detected encoded photons that had errors. The signal pulses correspond to pulses with excitation $S_{2}$, while the decoy pulses are those excited with $S_{1},S_{0}$. We show in the following that due to the truncated photon number basis of our SPS, only two different decoy states are enough for exact analysis in the framework of the decoy state protocol [11].
@@
On the other hand, due to the uncertainty of the photon-number in each pulse, the gain and error rate for a specific n-photon state, $\{Q_{n},e_{n}\}$, cannot be measured directly and have to be estimated or calculated. Out of this set, the single photon gain and error rate, $Q_{1},\,e_{1}$ which are the probability of a detected event to originate from a single photon pulse and the error rate for single photon pulses respectively are of particular importance. This is since, considering a possible PNS attack by Eve, only information encoded on the single photon pulses is secured.
Crucially, in our three-intensity DTB protocol, these parameters can be solved exactly since the probability distribution of the signal and decoy states, $\{P_{n}^{s}\},\{P_{n}^{d,i}\}$ is known, e.g. Fig. 1c. To relate the measured values $\{Q_{s},Q_{d,i},E_{s},E_{d,i}\}$ to the unknown set $\{Q_{n},e_{n}\}$, it is useful to define the n-photon yield, $Y_{n}$, as the conditional probability of a detection event given that Alice sent a n-photon state, and therefore $Q_{n}=Y_{n}P_{n}$. With these definitions, the following set of equations can be written [15]:
$Q_{s}$ $=\sum_{n=0}^{2}P_{n}^{s}Y_{n}\qquad E_{s}Q_{s}=\sum_{n=0}^{2}P_{n}^{s}Y_{n}e_{n}$ (6)
@@
Importantly, unlike the decoy state protocol implemented for WCS where $n\rightarrow\infty$, and therefore require infinite set of decoy states for an exact solution [11], here $n=0,1,2$ only. Thus, our equations have only six unknowns: $\{Y_{0},\,e_{0},\,Y_{1},\,e_{1},\,Y_{2},\,e_{2}\}$ suggesting that two decoy states are enough for an exact solution to these equations. This makes the DTB implementation particularly viable, whereas for WCS using two decoy states would only give a bounded approximation for the yield and errors [15].
The solution of these equations can then be used to estimate the minimum SKR after privacy amplification and error correction [11]:
$SKR\geq q\{-Q_{s}f(E_{s})H_{2}(E_{s})+Q_{1}\left[1-H_{2}(e_{1})\right]\}$ (8)
@@
Assuming $\eta_{C}\simeq 1$ for gCQD on a nanoantenna [36] and using the probability values- $P_{0},P_{1},P_{2}$, we can numerically calculate the SKR of such a source for different channel losses. For this, we use the channel model connecting the losses, yields, gains and the error-rates shown in [11] and in Supp. Sec. IV with the parameters values given in [41, 42], specifically, dark count probability of $P_{\text{dark}}=1.7\cdot{10}^{-6}$, probability to reach the wrong detector of $e_{d}=3.3\%$, Bob’s detection efficiency of $\eta_{\text{Bob}}=4.5\%$ and vacuum error-rate of $e_{0}=0.5$.
@@
Figure 3: QKD analysis with imperfect single photon sources. (a) Secure key rate of BB84 protocol as a function of channel loss for a perfect SPS (blue line), WCS with infinite decoy states and optimized intensities (WCS+Decoy, red), SPS1 with the decoy state on truncated basis protocol (SPS1+DTB, purple) and SPS2 with the heralded purification protocol (SPS2+HPP, black). The method and parameters of the calculations are found in Sec. III.1 and III.2. (b) The relative gain of the MCL, $\gamma$, as a function of $P_{1},P_{2}$ showing two advantageous regimes for SPS with either DTB or HPP separated by black dashed lines with their respective conditions. We also present our sources, SPS1 in the purification regime (black dot), SPS2 in the decoy state regime (purple square) and two, previously analyzed, non-classical sources [42, 43] marked with the orange square/dot.
@@
In Fig. 3(a) we show the calculated SKR under different channel losses for a standard WCS with decoy states (red) in the ideal, asymptotic case (infinite number of decoy states) with optimized intensity, compared to an SPS based on a gCQD coupled to a nanoantenna [36] (purple) using the above DTB protocol, with the realistically obtainable probabilities for such a device: $P_{1}^{s}=0.529$, $P_{2}^{s}=0.112$ [36] for the signal state. A clear improvement in the SKR and the MCL of our imperfect SPS over WCS is seen, with over $3$ dB MCL enhancement. We also show a comparison to a perfect single photon source ($P_{1}=1$) in black. Surprisingly, applying our truncated decoy-state protocol yields performance not far worse than a perfect SPS, even though our SPS is far from being ideal.
III.2 Heralded Purification Protocol- HPP

### v2

@@ Abstract
The original proposal of quantum key distribution (QKD) was based on ideal single photon sources, which 40 years later, are still challenging to develop. Therefore, the development of decoy state protocols using weak coherent states (WCS) from lasers, set the frontier in terms of secure key rates and distances. Here, we propose and experimentally demonstrate two simple-to-implement QKD protocols that allow practical, far from ideal sub-Poissonian photon sources to outperform state-of-the-art WCS. By engineering the photon statistics of a biexciton-exciton cascade in room temperature single photon sources based on giant colloidal quantum dots coupled to nanoantennas, we show that either a truncated decoy state protocol or a heralded purification protocol can be employed to achieve a significantly increased performance in terms of the maximal allowed channel loss for secure key creation, which can exceed even that of ideal WCS by more than 3dB. We then experimentally emulate a BB84 QKD using such a quantum dot source, verifying the superiority of our protocols over the best possible BB84 WCS performance. These protocols can be utilized efficiently on a host of various quantum emitters having controllable photon statistics with a finite photon-number basis, offering a practical approach to QKD without the hindering requirements on the single photon purity of the photon source.
@@ I Introduction
Quantum key distribution (QKD) stands as a prominent candidate for the post-quantum-computer optical communication, enabling a secure key-establishing protocol between two parties employing the quantum nature of light [1, 2, 3]. In theory, encoding the information for QKD onto pure single photon states is ultimately secured. In practice, it is limited by imperfections of the transmitter and the receiver, making security proofs a challenging task [4, 5], thus reducing the maximal secure key rate (SKR) and the maximal losses in the communication channel over which SKR is achievable [6, 7].
@@
Decoy state protocols introduce variations in the distribution of the emitted photons to better characterize the channel and obtain a tighter bound on the SKR [19, 20, 21, 22], allowing detection of PNS attacks [23, 24]. Yet, even with decoy state protocols, WCS are highly sensitive [25, 26] to the high probability of the vacuum state, due to the dependence between the different photon number emission probabilities, which overall limits the possible SKR [9]. The above solutions and inherent limitations of WCS for QKD set clear opportunities for improving the security and efficiency of QKD applications using realistic SPS systems, such as solid-state emitters [27, 28, 29, 30, 31, 32] and spontaneous parametric down conversion heralded sources [33, 34]. Yet, while efforts to develop a nearly ideal SPS are still ongoing by many groups [35, 36, 37, 38, 39], they still fall short as compared to the performance of QKD protocols based on WCS incorporating decoy states.
@@
Here we introduce a radically different, more realistic approach. By utilizing two new simple QKD protocols on compact, imperfect SPS sources, we significantly enhance the QKD performance, in terms of the MCL, as compared to even ideal, infinite decoy state protocols using WCS.
@@
We base our protocols on our recent demonstrations of room temperature operating SPS devices based on a giant colloidal nanocrystal quantum dot (gCQD) coupled to a hybrid nanoantenna and plasmonic nanocone [40, 41, 42, 43, 44]. These SPS devices displayed a highly enhanced rate of photon emission approaching the GHz range, together with near-unity collection efficiencies ($>90\%$) of the emitted photons. The photon emission from such gCQDs is based on the biexciton - exciton (BX-X) emission cascade under optical excitation [45], and has been shown to be well described by a truncated photon-number Fock space with $N=0,1$, and $2$ photons, with a negligible probability of $N>2$ photons [46].
@@
The two protocols are both based on our ability to vary and control the probabilities of emitting the different Fock states, $\{P_{0},P_{1},P_{2}\}$ over a wide range of values, by simply varying the optical excitation power. The first protocol we implement is a simple and realistic decoy state protocol for such a source. We show that with this new protocol, we can significantly exceed, by more than $3$ dB, the MCL value of an ideal decoy state protocol using a WCS source. The second protocol uses an alternative approach based on heralded photon purification under saturated excitation. We show that a similar performance enhancement can also be achieved. Later, we experimentally verify that the probability for $N>2$ ($P_{N>2}$) is negligibly small, justifying our protocols assumptions. Finally, we experimentally fully emulate a free-space BB84 QKD using a bare gCQD source without an antenna. The results agree well with the models, and demonstrate the superiority of our protocols over the best possible BB84 WCS performance even for such a simple imperfect SPS source. This shows that even existing SPS, which are far from being ideal, can outperform the current state-of-the-art protocols based on WCS sources.
@@
Figure 1: (a) Schematic of the BX-X cascade in an optically excited gCQD, using a non-resonant pulsed excitation with a normalized intensity $S=\frac{I}{I_{S}}$. Inset is a sketch of a typical gCQD, consisting of a CdSe core and a CdS shell. (b) Normalized emission counts (blue open circles) as a function of $S$, displaying a saturation behaviour. The dashed black line is a fit to a model, detailed in Section II. The orange circles show the experimentally extracted $g^{(2)}(0)$ for the different powers. (c) Photon probabilities as a function of $S$. The dashed black lines mark the intensities chosen for the signal $S_{2}$, and the two decoy states $S_{0}$, $S_{1}$, for the decoy state on a truncated basis (DTB) protocol, and the intensity $S_{3}$, chosen for the heralded purification (HP) protocol.
@@ II Optical control of photon statistics from the BX-X emission cascade in a gCQD
Fig. 1(a) presents a schematic sketch of the BX-X cascade in a gCQD (see Appendix A) following a non-resonant pulse excitation with power $I$. Modelling this process, a BX state |$\ket{XX}$ is excited by the absorption of two pump photons. This process happens with a probability:
@@
Since $P_{XX},P_{X}$ depend on the excitation power, the probabilities to emit two, one or zero photons following excitation become excitation power dependent, allowing an external control over the emitted photon statistics as is required for decoy states protocols. Such a saturable behavior of the excitonic population allows us to define a relative excitation power $S=I/I_{S}$, where $I_{S}$ is the saturation power, for which the detected count rate (which is linearly proportional to the average emission photon number) reaches $90\%$ (a convenient arbitrary value) of the maximum value (which asymptotically occurs at $I\rightarrow\infty$).
@@
In Fig. 1(b), we experimentally demonstrate photon statistics control by varying the excitation power from a single bare gCQD. As $S$ increases, the collected photon emission intensity increases sublinearly and saturates at $S\gtrsim 1$. A good fit of the theoretical model, based on the absorption probabilities given in Eqs. 1, 2 and on the subsequent emission probabilities $QY_{X},QY_{XX}$ is shown (see Appendix B for details). This fit confirms the basic assumptions given above for the BX-X cascade process.
@@
At the same time, the measured second-order correlation function, $g^{(2)}(0)$, increases in a similar manner. This is understood in the following way: at low powers, most emission events are empty as $P_{0}$ is the dominant term. As $S$ increases, $P_{1}$ increases linearly (Eq. 2), while $P_{2}$ increases quadratically and thus is negligible at $S\ll 1$. Since $g^{(2)}(0)$ is an increasing function of $P_{2}/(\sum_{i=0,1,2}{P_{i}})$, a saturable behavior of $g^{(2)}(0)$ with increasing $S$ is also expected.
@@
Importantly, as we also show experimentally (see Appendix D), multi-excitonic ($N>2$) emissions from the gCQD can be neglected due to their low emission probability, thus $P_{N>2}=0$.
@@
By measuring both $g^{(2)}(0)(S)$ and the count-rate $C(S)$, we can extract the probability distribution precisely, as is detailed in Appendix C.
@@
In Fig. 1(c), we present the extracted photon emission probabilities $\{P_{0},P_{1},P_{2}\}(S)$. As expected, the probability of emitting one or two photons increases with excitation power, but at a different rate, allowing optical control of the emitted photon statistics with only varying $S$.
@@
Figure 2: Concept of two protocols based on a high rate, high collection efficiency SPS with imperfect purity: (a) Experimental setup and sketch of the BB84+DTB protocol, with Alice adjusting the excitation power on the SPS between $S_{0}=0$, $S_{1}$ and $S_{2}$ to control the emitted photon statistics for the 2-decoy protocol, followed by collection optics with efficiency $\eta_{C}$ and a standard BB84 encoding unit. (b) Experimental setup and sketch of the BB84+HP photon purification protocol. A single excitation power $S_{3}>1$, near the gCQD saturation, is used to excite the SPS followed by collection optics with efficiency $\eta_{C}$, a beam-splitter (BS), a high efficiency single photon detector with detection efficiency $\eta_{D}$, and a standard BB84 encoding unit. Here, only same pulse detections at Alice and Bob are used for the secure key. An algorithmic representation of the two protocols is given in Appendix H.
@@ III.1 Decoy State on a Truncated Fock Basis - DTB
In Fig. 2(a), we present the conceptual experimental setup of the decoy state protocol. Here, an SPS with a near unity photon collection efficiency, $\eta_{c}\simeq 1$ [43] (titled as SPS1), is used for implementing a BB84 protocol, but where Alice controls the excitation powers, $S_{0}=0$, $S_{1}$ and $S_{2}$. This in turn modifies the photon number statistics, thus allowing for the implementation of decoy and signal states [11, 15].
To establish the DTB protocol, we define the gain of the signal (decoyi) state, $Q_{s}$ ($Q_{d,i}$), as the fraction of encoded photon pulses sent by Alice and detected by Bob (Eq. 11). We also define the error rate of the signal (decoyi), $E_{s}$ ($E_{d,i}$), as the fraction of detected encoded photons that had errors (Eq. 12). The signal pulses correspond to pulses with excitation $S_{2}$, while the decoy pulses are those excited with $S_{1},S_{0}$. We show in the following that due to the truncated photon number basis of our SPS, only two different decoy states are enough for exact analysis in the framework of the decoy state protocol [11].
@@
On the other hand, due to the uncertainty of the photon-number in each pulse, the gain and error rate for a specific n-photon state, $\{Q_{n},e_{n}\}$, cannot be measured directly and have to be estimated or calculated. Out of this set, the single photon gain ($Q_{1}$) and error rate ($e_{1}$), which are the probability of a detected event (Bob) to originate from a single photon pulse (Alice) and the error rate for single photon pulses respectively are of particular importance. This is since, considering a possible PNS attack by Eve, only information encoded on the single photon pulses is secured.
Crucially, in our three-intensity DTB protocol, these parameters can be solved exactly since the probability distribution of the signal and decoy states, $\{P_{n}^{s}\},\{P_{n}^{d,i}\}$ is known, e.g. Fig. 1(c). To relate the measured values $\{Q_{s},Q_{d,i},E_{s},E_{d,i}\}$ to the unknown set $\{Q_{n},e_{n}\}$, it is useful to define the n-photon yield, $Y_{n}$, as the conditional probability of a detection event given that Alice sent a n-photon state, and therefore $Q_{n}=Y_{n}P_{n}$. With these definitions, the following set of equations can be written [15]:
$Q_{s}$ $=\sum_{n=0}^{2}P_{n}^{s}Y_{n}\qquad E_{s}Q_{s}=\sum_{n=0}^{2}P_{n}^{s}Y_{n}e_{n}$
@@
Importantly, unlike the decoy state protocol implemented for WCS where $n\rightarrow\infty$, which requires an infinite set of decoy states for an exact solution [11], here $n=0,1,2$ only, due to the truncated sub-Poissonian nature of the gCQD photon emission ($N>2$ photon emissions are neglected). Thus, our equations have only six unknowns: $\{Y_{0},\,e_{0},\,Y_{1},\,e_{1},\,Y_{2},\,e_{2}\}$, suggesting that two decoy states are enough for an exact solution to these equations. This makes the DTB implementation particularly viable, whereas for WCS using two decoy states would only give a bounded approximation for the yield and errors [15].
The solution of these equations can then be used to estimate the minimum SKR after privacy amplification and error correction, defined as $R_{1}$ for the DTB protocol [11]:
$R_{1}\geq q\{-Q_{s}f(E_{s})H_{2}(E_{s})+Q_{1}\left[1-H_{2}(e_{1})\right]\}$ (7)
@@
Assuming $\eta_{C}\simeq 1$ for a gCQD coupled to a nanoantenna [43] and using the $P_{0},P_{1},P_{2}$ probability values above, we numerically calculate the SKR of such a source for different channel losses. To do this, we use the channel model connecting the losses, yields, gains and the error rates shown in [11] and in Appendix E with the parameters values (from [48, 49, 41, 43]) presented in Table 1, Appendix G. Specifically, we set the probability to reach the wrong detector to $e_{d}=3.3\%$ and Bob’s detection efficiency to $\eta_{\text{Bob}}=4.5\%$.
@@
Figure 3: QKD analysis with imperfect single photon sources. (a) Secure key rate of the BB84 protocol as a function of channel loss for a perfect SPS (blue line), WCS with infinite decoy states and optimized intensities (WCS+Inf. Decoy, red), SPS1 with the decoy state on a truncated Fock basis protocol (SPS1+DTB, purple), SPS2 with the heralded purification protocol (SPS2+HP, black), and a typical gCQD (bare gCQD, green), with a visual representation of $\gamma$, the relative gain. The method and parameters of the calculations are found in Table 1 in Appendix G, Sec. III.1, and Sec. III.2. (b) The relative gain of the MCL, $\gamma$, as a function of $P_{1},P_{2}$ showing two advantageous regimes for SPS with either DTB or HP separated by black dashed lines with their respective conditions. We also present our sources: a bare single gCQD experimentally measured for the DTB protocol (purple circle) and the HP scheme (black circle) at different intensities (hence different photon statistics), SPS1 in the purification regime (purple square), SPS2 in the decoy state regime (black square), and two, previously analyzed, non-classical sources [49, 50] marked with an orange square/dot.
@@
In Fig. 3(a) we show the calculated SKR under different channel losses for a standard WCS with decoy states in the ideal, simulated, asymptotic case (infinite number of decoy states) with optimized Poissonian distribution parameters, as given by Ref. [15], setting the detection efficiency and error probability as previously explained. This was compared to an SPS based on a gCQD coupled to a nanoantenna (SPS1) [43] using the above DTB protocol with the realistically obtainable probabilities of the signal state for such a device (see Appendix G). A clear improvement in the SKR and the MCL of our imperfect SPS over WCS is seen, with over $3$ dB MCL enhancement, and also an improvement over the best existing cryogenic state-of-the-art SPS [31, 28, 32] or comparable cryogenic, fiber-coupled results [51]. We also show a comparison to a perfect single photon source ($P_{1}=1$). Surprisingly, applying our truncated decoy-state protocol yields performance not far worse than a perfect SPS, even though our SPS, operating under ambient conditions, is far from being ideal.
III.2 Heralded Purification - HP